OpenStudy (anonymous):

Determine if the summation (to be posted below) is convergent or divergent through the alternating series test.

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}(-1)^n(\sqrt{n+1}-\sqrt{n})\]

OpenStudy (anonymous):

when they state the alternating series test in my textbook they have the -1 to the power of n-1 not n so I have to change it to \[\sum_{n=1}^{\infty}(-1)^{n-1}(\sqrt{n}-\sqrt{n+1})\] right?

OpenStudy (anonymous):

but thats increasing so how am I suppose to use AST?

OpenStudy (primeralph):

Either way, it's decreasing absolutely.

OpenStudy (anonymous):

heres the thing the instructor said to use the concepts they introduced in that section to answer these questions. All they introduced this section was the AST

OpenStudy (primeralph):

Just check if \[|\lim_{n \rightarrow \infty} \frac{ A _{n+1} }{ A _{n} } |<1\]

OpenStudy (anonymous):

I know the ratio test but unfortunately I can't use that yet. Its introduced in a later section in the textbook

OpenStudy (primeralph):